16x²+ 9y²+64x-54y+1=0 center
Found 2 solutions by ikleyn, Theo:
Answer by ikleyn(52802)

.
Learn how to communicate with other people,  and with the tutors in this forum,  in particular.


Answer by Theo(13342)

looks like the equation of an ellipse.

assuming that's what it is, then transform the equation into standard form of an ellipse.

start with 16x^2 + 9y^2 + 64x - 54y + 1 = 0

subtract 1 from both sides to get 16x^2 + 9y^2 + 64x - 54y = -1

group the x's and the y's together to get (16x^2 + 64x) + (9y^2 - 54y) = -1

factor out the coefficients so that the grouped variables start with a coefficient of 1.

you will get 16(x^2 + 4x) + 9(y^2 - 6y) = -1

complete the squares on (x^2 + 4x) to get (x+2)^2 - 4
complete the squares on (y^2 - 6y) to get (y-3)^2 - 9

equation becomes 16 * ((x+2)^2 - 4) + 9 * ((y-3)^2 - 9) = -1

simplify to get 16 * (x+2)^2 - 64 + 9 * (y-3)^2 - 81 = -1

add 64 and 81 to both sides of the equation to get:

16 * (x+2)^2 + 9 * (y-3)^2 = 144

divide both sides of the equation by 144 to get:

16 * (x+2)^2 / 144 + 9 * (y-3)^2 / 144 = 1

this can be written as 16/144 * (x+2)^2 + 9/144 * (y-3)^2 = 1

simplify so that the numerator in the fractions is 1 to get (x+2)^2 / 9 + (y-3)^2 / 16 = 1

this is now in standard form of (x-h)^2 / b^2 + (y-k)^2 / a^2 = 1

the designation of the letter a always goes where the largest denominator is and the designation of the letter b always goes where the smallest denominator is.

that's why the b^2 went under the (x-h)^2 and the a^2 went under the (y-k)^2.

in your equation, you have:

a^2 = 16 which makes a = 4
b^2 = 9 which makes b = 3
center of the ellipse is (h,k) which make the center (-2,3).

the major axis of the ellipse is the axis that is the longest.

a is the distance along the major axis from the center of the ellipse to the vertex of the ellipse.

b is the distance along the minor axis from the center of the ellipse to the co-vertex of the ellipse.

c is the distance along the major axis from the center of the ellipse to the the focus of the ellipse.

c is calculated using the formula c^2 = a^2 - b^2.

in your equation, c^2 is therefore equal to 9 - 4 = 5.

this makes c = sqrt(5).

your important values to the ellipse are now:

a = 4
b = 3
c = sqrt(5)
center = (-2,3)

the graph of your ellipse looks like this:

$$$

here's a reference that should help you understand what's going on.

http://www.purplemath.com/modules/ellipse.htm

here's my diagram of the main parts of the ellipse that should help you identify what's what on the ellipse.

$$$